Evaluate the following functions in differential operator form.

  1. D2((x+1)2/x-1) Answer: 8/(x-1)3
  2. D(ln(x2+2x+1) Answer: 2/x-1
1
Expert's answer
2021-12-06T15:02:07-0500

Q1.

Let y = "\frac{(x+1)²}{x-1}"

So y = "\frac{(x-1)²+4x}{x-1}= \frac{(x-1)²+4(x-1)+4}{x-1}"

=> y = (x-1) +4 + "\frac{4}{x-1}"

=> y = x + 3 + "\frac{4}{x-1}"

So "\frac{dy}{dx}= 1+0-\frac{4}{(x-1)²}"

Differentiating again

"\frac{d²y}{dx²}= 0-\frac{4*(-2)}{(x-1)³}= \frac{8}{(x-1)³}"

Therefore D²("\frac{(x+1)²}{x-1}" ) = "\frac{8}{(x-1)³}"

Q2. D(ln(x²+2x+1))

= D(ln(x+1)²)

= D(2ln(x+1)) since ln(A)n = n ln(A)

= 2D(ln(x+1))

= "2\frac{d}{dx}(ln(x+1))"

= "\frac{2}{x+1}"


Note: Given answer of question number 2 is wrong.